# SupplementaryMat

### v1

@@ Abstract
We study spatiotemporal chaos in dense active suspensions using a generalized hydrodynamic model. As the activity is increased, we observe a clear structural transition marked by the formation of intense, spatially organized vortices with the appearance of giant number fluctuations. This transition is accompanied by the emergence of local polar ordering, an increase in velocity correlations, and a shift in the velocity distribution from unimodal to bimodal. Beyond activity, reducing the instability timescale leads to similar structural changes in the flow. This suggests that both activity and instability dynamics can serve as control parameters for pattern formation. Our findings reveal how local ordering and dynamical instability interact to generate large-scale structures in active systems. Furthermore, based on the energy budget, we define an order parameter that quantifies the structural changes in the phase space of activity and instability time scales.
@@
Introduction: The collective behavior of active particles gives rise to complex and intriguing phenomena, including large-scale coherent motion in flocks, mesoscale turbulence in bacterial suspension, and defect-driven chaos in active liquid crystals alert2022active ; marchetti2013hydrodynamics ; toner2005hydrodynamics ; giomi2013defect ; james2021emergence ; ngo2014large ; guillamat2016control ; wensink2012meso ; sokolov2007concentration ; martinez2020combined ; guillamat2017taming ; wei2024scaling ; rorai2022coexistence . The spatiotemporal structures observed in dense suspensions of active matter qualitatively resemble those found in passive turbulence systemswensink2012meso ; alert2022active ; martinez2020combined ; wolgemuth2008collective ; dombrowski2004self ; sokolov2007concentration ; doostmohammadi2019coherent . Extensive studies have been conducted to investigate whether active turbulence shares similarities with or deviates from inertial turbulence. These investigations often focus on identifying universal features, such as power-law scaling and intermittency, which are hallmarks of turbulent systems cp2020friction ; martinez2021scaling ; mukherjee2023intermittency ; kiran2023irreversibility ; Pandit_2025 . Transitions in the flow characteristics are observed at critical values of the activity parameter, characterized by increased intermittency and deviations from Gaussian statistics mukherjee2023intermittency ; kiran2025onset . Until recently, there was no consensus on the existence of universal features in active turbulence. A recent study led by Ray mukherjee2023intermittency reported the emergence of universal scaling behavior in the energy spectrum beyond a critical activity threshold, highlighting the intricate interplay between active forcing and the system’s underlying field dynamics.
@@
However, there is also evidence pointing to structural changes in the vorticity field itself with increasing activity ramaswamy2016activity ; linkmann2019phase ; linkmann2020condensate ; mukherjee2021anomalous ; puggioni2022giant . This highlights the need for a deeper understanding of how these emergent structures contribute to the observed universal and nonuniversal features of active turbulence. Gaining insight into these structural changes and flow dynamics is important for explaining various processes like nutrient mixing and molecular transport in biological and ecological systems. sokolov2009reduction ; kurtuldu2011enhancement ; dombrowski2004self ; hernandez2005transport .
@@
Figure 1: The vorticity field ($\omega$) is shown for two different activity levels: (a) $\alpha=-1$, and (b) $\alpha=-9$. The scale bar in both panels corresponds to $25\Lambda$. Panels (c) and (d) present zoomed-in sections of the respective vorticity fields, with magnified regions chosen to be of the same order as the scale bar. The vorticity snapshot at higher activity displays spatial inhomogeneities, characterized by intense vortex regions coexisting with areas of negligible vorticity. The color bar is normalized by the maximum vorticity $\omega_{0}$ of the respective $\alpha$.
@@
In this work, we employ hydrodynamic models developed to describe the complex, chaotic flows in dense, quasi-two-dimensional bacterial suspensions wensink2012meso ; mukherjee2023intermittency ; kiran2025onset . These models have also been used to study turbulence-like behavior in dense suspensions of Bacillus subtilis alert2022active ; Pandit_2025 . We focus on the emergence and evolution of spatiotemporal structural patterns as a function of an activity parameter. Increasing activity leads to a strikingly uneven distribution of intense vorticity centers in space and time; see, for example, Fig 1. Tracking the number of vortex centers reveals the emergence of giant number fluctuations, hinting at an underlying structural transition in the flow. To probe the underlying origin of this transition, we examine the spatial velocity correlations and observe a marked increase in correlation length, along with the onset of local polar ordering. Interestingly, these locally ordered regions do not take over the whole system; instead, they coexist with disordered, chaotic zones. Similar structures have been observed experimentally wei2024scaling . Furthermore, we find that tuning the linear instability growth timescale—introduced to model the intrinsic instabilities in dense active suspensions—can alter the system’s phase even at fixed activity levels, effectively shifting the critical activity threshold for structural transitions.
Generalized Hydrodynamic model: We consider a coarse-grained hydrodynamic model to describe the collective dynamics of a two-dimensional dense bacterial suspension wensink2012meso :
[…]
@@
where $\bm{u}(\bm{x},t)$ is the coarse-grained velocity field. The divergence-free condition enforces the incompressibility of $\bm{u}$. $p(\bm{x},t)$ is the pressure field. The coefficient $\lambda_{0}$ encodes the effects of different swimmer types: $\lambda_{0}>1$ corresponds to pusher-type swimmers, while $\lambda_{0}<1$ corresponds to puller-type swimmers. The Toner-Tu coefficients $(\alpha,\beta)$ describe the polar ordering in self-propelled particle systems toner2005hydrodynamics : for values of $\alpha<0(>0)$ the term $\alpha\boldmath{u}$ injects (dissipates) energy and $\beta>0$ dissipates energy while ensuring numerical stability. Together, they define a characteristic velocity scale, $v=\sqrt{-\alpha/\beta}$, for polar-ordered motion. The ($\Gamma_{0},\Gamma_{2}$) term has the structure of the Swift-Hohenberg equation swift1977hydrodynamic . With $\Gamma_{0}<0$ and $\Gamma_{2}>0$, this term inherently drives the system towards a linear instability at length scales, $\Lambda=2\pi\sqrt{\frac{2\Gamma_{2}}{\Gamma_{0}}}$. It is the interplay between these terms that leads to chaotic, self-sustaining, turbulent-like flow patterns.
@@
Simulation details: To numerically solve Eqs. (1), we employed a de-aliased pseudo-spectral algorithm with periodic boundary conditions in space. We perform two-dimensional simulations in a square domain of size $160\times 160$ and grid resolution $4096\times 4096$, ensuring sufficient resolution for capturing fine-scale features. Time integration was performed using the second-order integrating factor method (IFRK2) with a time step of $\Delta t=2\times 10^{-4}$. The simulations were initialized with random initial conditions and in all simulation cases, we set the parameters fixed unless otherwise stated, $\lambda_{0}=3.5$, $\beta=0.5$, $\Gamma_{2}=9\times 10^{-4}$, $\Gamma_{0}=4.5\times 10^{-2}$. These values are chosen based on previous studies that align with experimentally accessible regimes kiran2023irreversibility . For more details, see section S1. To investigate the role of activity, we systematically vary the parameter $\alpha$ between -9 to -1.
Figure 2: (a) Probability distribution functions (PDF) of the normalized vortex count (n/$(\rm n/\bar{n})$ in subdomains of fixed linear size $\ell$ for different activity values. As $|\alpha|$ increases the distributions broaden and their maxima shift from $\rm n/\bar{n}\simeq 1$ toward lower values, indicating that typical subdomains contain fewer vortex centers, while strong fluctuations become more likely. (b) Standard deviation $\Delta n$ of the number of vortex centers within subregions is plotted against the corresponding mean $\bar{\rm n}$ as the subregion size $\ell$ increases. The scaling behavior $\Delta\rm n\sim\bar{n}^{\delta}$ reveals the presence of giant number fluctuations for $\alpha<\alpha_{c}$. Inset: the scaling exponent $\delta$ is plotted as a function of the activity parameter $\alpha$, showing that fluctuations become increasingly pronounced (larger $\delta$) with higher activity.
Figure 3: (a) Vorticity field for the entire domain for $\alpha=-9$, with the scale bar indicating a length scale of $25\Lambda$ (b) and (c) show the velocity streamlines overlaid on the velocity magnitude for a domain of size $25\Lambda$. (d) The normalized velocity correlation length $\zeta$ versus $\alpha$. Inset: The spatial velocity correlation function, defined as $C_{v}(r)=\langle\mathbf{u}(\mathbf{r}_{0})\cdot\mathbf{u}(\mathbf{r}_{0}+\mathbf{r})\rangle$ is shown as a function of distance $r$ for increasing values of $|\alpha|$. The velocity correlation length $\zeta$ is extracted by fitting the decay of $C_{v}(r)$ to an exponential form $C_{v}(r)\sim e^{-r/\zeta}$. For $\alpha>\alpha_{c}$, the normalized correlation length $\xi/\Lambda$ remains close to one, whereas it increases significantly for $\alpha<\alpha_{c}$, indicating the emergence of long-range velocity correlations in the high-activity regime.
@@
Spatial distribution of vortex centers: We show in Fig. 1 (a) and (b) the plots of vorticity field $\omega(\mathbf{x},t)=\nabla\times\mathbf{u}(\mathbf{x},t)$ for different values of the activity parameter $\alpha$. The field $\omega$ undergoes a structural transformation as $\alpha$ decreases, with large-scale spatial inhomogeneities becoming more prominent at lower $\alpha$ or, equivalently, higher activity. These inhomogeneities are apparent from the long-range correlations in the vorticity field as seen in Figs. 1(b) and  1(d). In active matter systems, such large-scale spatial inhomogeneities are indicative of giant number fluctuations where the standard deviation ($\Delta\rm n$) of the number of active particles ($\rm n$) grows anomalously large and scales as $\Delta\rm n\sim\rm n^{\delta}$, with the exponent $\delta$ reaching values as high as 1 in two dimensions chate2020dry ; narayan2007long ; fily2012athermal . In the asymptotic limit of $\rm n$, $\delta=0.5$ for systems in equilibrium and away from continuous phase transitions. In analogy, in active turbulence, we find giant number fluctuations of the vortex centers. To quantify the fluctuations in the number of vortices, we compute the average number of vortices $\bar{\rm n}$ and the corresponding variance, $\Delta\rm n$ within a subregion of size $l^{2}$. In Fig. 2(a), we plot the probability distribution (PDF) of $\rm n$ around its mean for different $\alpha$, in a given subregion. We note that for decreasing $\alpha$, the $P(\rm n)$ broadens, indicating stronger fluctuations in the number of vortex centers. In Fig. 2(b) we plot $\Delta\rm n$ versus $\bar{\rm n}$ for different activities and find that $\Delta\rm n\sim\bar{\rm n}^{\delta}$, with $\delta\simeq 0.5$ for $\alpha\geq-5$ and $\delta>0.5$ for $\alpha<-5$. The values of $\delta$ approach 1 for very large $|\alpha|$, see inset of Fig. 2(b). Notably, the transition of $\delta$ at a critical value of the activity parameter, $\alpha_{c}=-5$, matches with the previously identified critical activity by the group of Ray mukherjee2021anomalous ; mukherjee2023intermittency through the onset of universality in energy spectra, increased intermittency, and other hallmarks of turbulent transitions. Note that with increasing activity, although the vorticity at individual vortex centers increases, the total number of such centers decreases S2(b).
@@
The methodology for identifying vortex centers is provided in section S2. While the fluctuations in number reflect statistical deviations, a more structural perspective on the spatial inhomogeneity of vortex centers is offered by Voronoi tessellation. The results of this analysis are consistent with the emergence of giant number fluctuations at high activity levels (see section S3).
@@
Emergence of local ordering : This statistical signature of inhomogeneity naturally raises the question: what changes in the underlying flow field are responsible for this transition? To address this, we closely examine the structure of the vorticity and velocity fields. For $\alpha<\alpha_{c}$, there are a few large vortical regions of the size ${}\sim 25\Lambda$ characterized by low or nearly vanishing vorticity (see Fig. 1(d)) interspersed with vortical regions of the size $\mathcal{O}(\Lambda)$. In Fig 3(b), we plot the velocity streamlines corresponding to these large vortical regions of vanishing vorticity and note that they are highly aligned, indicating local ordering. This contrasts sharply with other regions of the same field Fig 3(c), where the flow remains disordered and lacks directional alignment. Furthermore, in Fig 3(d) we plot velocity correlation $C_{v}(r)=\langle\mathbf{u}(\mathbf{r}_{0})\cdot\mathbf{u}(\mathbf{r}_{0}+\mathbf{r})\rangle$ versus $r$ and see that there is a sharp transition followed by a monotonic increase in correlation length, $\zeta$, for $\alpha<\alpha_{c}$. For sufficiently negative $\alpha$ (e.g., $\alpha\lesssim-6$), the spatial velocity correlation length reaches up to $\mathcal{O}(10\Lambda)$ and for $\alpha=-9$ we find $\zeta\simeq 25\Lambda$, consistent with the lengths of the large vortical regions seen in Fig. 1 (b). For $\alpha>\alpha_{c}$, we note from Fig. 1 (a) and (c), the largest vortical structures are $\sim\mathcal{O}(\Lambda)$ and velocity streamlines are chaotic throughout domain, see Fig. S2(a). Signatures of local ordering can also be seen in the PDF of the velocity field: we plot in Fig. 4(a) the PDF of a component of the velocity field for different activities and observe that the PDF transitions from a unimodal distribution centered at $\langle u\rangle=0$ for $\alpha>\alpha_{c}$ to a bimodal for $\alpha<\alpha_{c}$. The two peaks in the bimodal state progressively approach ±$\pm\sqrt{\frac{\alpha}{\beta}}$, consistent with the predictions of the Toner-Tu model for self-propelled active systems. This highlights competition between active alignment-induced ordering Toner-Tu terms and spatial modulation introduced by higher-order gradient terms.
@@
The global phase behavior is not dictated by a single control parameter but instead emerges from a balance between multiple energetically distinct contributions. The first is the energy associated with activity-driven motion, quantified as $E_{\rm active}=\frac{v^{2}}{2}.$ The second arises from the destabilizing active stresses introduced by the generalized Swift–Hohenberg-like term, which generates intrinsic instabilities and injects energy into the system at intermediate length scales. The energy associated with this contribution is given by $E_{\Lambda}=\frac{|\Gamma_{0}^{3}|}{\Gamma_{2}}.$ The third component is the kinetic energy of the flow, $E_{\rm kin}(\equiv\int_{\mathbf{x}\in L^{2}}\frac{1}{2}\mathbf{u}\cdot\mathbf{u}d\mathbf{x}$). This energetic framework allows us to define an effective order parameter, $\phi$, that captures the interplay between these different forms of energy input and transfer.
@@
We hypothesize that the relative magnitudes of $E_{\rm active}$, $E_{\Lambda}$, and $E_{\rm kin}$ determine the qualitative state of the system — whether it exhibits local ordering identified by long correlation lengths or remains in a disordered phase. For $\alpha<\alpha_{c}$, the order parameter $\phi$ becomes positive, indicating that the active energy $E_{active}$ dominates over other contributions. This leads to a breakdown of homogeneity and the emergence of local ordering. In contrast, $\phi<0$ corresponds to a homogeneous, disordered state where active energy is insufficient to induce such organization, Fig. 4(b).
Figure 4: (a) Probability distribution function of the a velocity component $u_{x}$ for different values activity $\alpha$. A transition from an unimodal to a bimodal distribution is observed as $\alpha$ decreases beyond the critical point $\alpha_{c}$. (b) The order parameter $\phi=E_{\rm active}-E_{\Lambda}-E_{\rm kin},$ is plotted as a function of $\alpha$. $\phi$ is positive below $\alpha_{c}$ and becomes negative above it, reflecting the onset of the transition.

### v2

# Emergence of Local Ordering and Mesoscale Giant Number Fluctuations in Active Turbulence

@@ Abstract
We study spatiotemporal chaos in two-dimensional dense active suspensions using a generalized hydrodynamic model. Increasing activity induces a structural transition marked by the formation of intense vortices and giant number fluctuations at the mesoscale. The flow self-organizes into locally polar-ordered regions coexisting with chaotic domains, producing a bimodal velocity distribution and enhanced correlations. This mixed-state morphology underlies the universal statistical behavior observed beyond a critical activity threshold. Reducing the instability timescale yields similar transitions, showing that both activity and instability act as control parameters for pattern formation. An energy-based order parameter derived from the system’s budget quantifies and unifies these structural transitions across the phase space of activity and instability timescales.
@@
Introduction—The collective behavior of active particles gives rise to complex and intriguing phenomena, including large-scale coherent motion in flocks, mesoscale turbulence in bacterial suspension, and defect-driven chaos in active liquid crystals [1, 22, 44, 43, 11, 15, 29, 12, 48, 40, 23, 13, 46, 38]. The spatiotemporal structures observed in dense suspensions of active matter qualitatively resemble those found in passive turbulence systems[48, 1, 23, 49, 6, 40, 7]. Extensive studies have been conducted to investigate whether active turbulence shares similarities with or deviates from inertial turbulence. These investigations often focus on identifying universal features, such as power-law scaling and intermittency, which are hallmarks of turbulent systems [4, 24, 27, 16, 32]. Interestingly, the statistical properties at high activity resemble those of inertial turbulence—exhibiting scaling and intermittency: recent studies demonstrated the emergence of universal scaling and intermittency in the energy spectrum beyond a critical activity threshold [27, 17].
While these statistical features closely resemble those of inertial turbulence, their physical origin in active systems remains unclear. In particular, it is not known whether the universal scaling observed at high activity arises from fully chaotic dynamics or from a more complex coexistence of ordered and disordered regions.
@@
In this work, we demonstrate that beyond a critical activity threshold, the flow field self-organizes into locally ordered domains coexisting with chaotic regions, and that this mixed-state structure underlies the universal behavior observed in active turbulence. By systematically varying the activity and instability timescale, we show that both serve as control parameters governing the onset of local ordering and the associated structural transitions. However, there is also evidence pointing to structural changes in the vorticity field itself with increasing activity [36, 19, 20, 26, 34, 10, 2]. This highlights the need for a deeper understanding of how these emergent structures contribute to the observed universal and nonuniversal features of active turbulence. Gaining insight into these structural changes and flow dynamics is important for explaining various processes like nutrient mixing and molecular transport in biological and ecological systems [41, 18, 6, 14].
@@
Figure 1: The vorticity field ($\omega$) is shown for two different activity levels: (a I) $\alpha=-1$, and (a II) $\alpha=-9$ at $t=85\tau_{\Gamma}$ with $\tau_{\Gamma}=0.4$. The scale bar in both panels corresponds to $25\Lambda$. Panels (a III) and (a IV) present zoomed-in sections of the respective vorticity fields, with magnified regions chosen to be of the same order as the scale bar. The vorticity snapshot at higher activity displays spatial inhomogeneities, characterized by intense vortex regions coexisting with areas of negligible vorticity. The color bar is normalized by the maximum vorticity $\omega_{0}$ of the respective $\alpha$. (b I) and (b II) display the velocity streamlines superimposed on the velocity magnitude within a subdomain of size $25\Lambda$, corresponding to the regions marked by the red and green boxes in (a II). (c) Standard deviation $\Delta n$ of the number of vortex centers within subregions is plotted against the corresponding mean $\bar{n}$ as the subregion size $\ell$ increases. The scaling behavior $\Delta n\sim\bar{n}^{\delta}$ reveals the presence of giant number fluctuations for $\alpha<\alpha_{c}$. Inset: the scaling exponent $\delta$ is plotted as a function of the activity parameter $\alpha$, showing that fluctuations become increasingly pronounced (larger $\delta$) with higher activity.
@@
We employ hydrodynamic models developed to describe the complex, chaotic flows in dense, quasi-two-dimensional bacterial suspensions [48, 27, 17]. These models have also been used to study turbulence-like behavior in dense suspensions of Bacillus subtilis [1, 32]. While previous studies have primarily focused on statistical signatures such as energy spectra or intermittency, the energetic underpinnings of these transitions remain poorly understood. We introduce an energy-based order parameter that provides a unifying framework to distinguish ordered and disordered regimes in active turbulence, linking structural organization directly to the system’s underlying energy balance and examining how the flow field reorganizes as activity and instability compete. Increasing activity leads to a strikingly uneven distribution of intense vorticity centers in space and time (Fig 1(a)). Tracking the number of vortex centers reveals emergence of giant number fluctuations at the mesoscale, i.e., at length scales larger than the characteristic vortex size but smaller than the system size, hinting at an underlying structural transition in the flow. To probe the underlying origin of this transition, we examine the spatial velocity correlations and observe a marked increase in correlation length, along with the onset of local polar ordering. Interestingly, these locally ordered regions do not take over the whole system; instead, they coexist with disordered, chaotic zones. Furthermore, we find that tuning the timescale of the linear instability, introduced to model intrinsic instabilities in dense active suspensions, can alter the system’s phase even at fixed activity levels. This effectively shifts the critical activity threshold for structural transitions.
Generalized hydrodynamic model— Following the hydrodynamic frameworks proposed for dense bacterial suspensions [48], we describe the system using an incompressible velocity field $\mathbf{u}(\mathbf{x},t)$ that evolves according to :
$\begin{aligned} \frac{\partial\mathbf{u}}{\partial t}+\lambda_{0}\mathbf{u}\cdot\nabla\mathbf{u}&=-\nabla p-\left(\alpha+\beta|u|^{2}\right)\mathbf{u}+\Gamma_{0}\nabla^{2}\mathbf{u}-\Gamma_{2}\nabla^{4}\mathbf{u}\\ \nabla\cdot\mathbf{u}&=0\end{aligned}$
@@
$p(\bm{x},t)$ is the pressure field. The coefficient $\lambda_{0}$ encodes the effects of different swimmer types: $\lambda_{0}>1$ corresponds to pusher-type swimmers, while $\lambda_{0}<1$ corresponds to puller-type swimmers. The Toner-Tu coefficients $(\alpha,\beta)$ describe the polar ordering in self-propelled particle systems [44, 43]: for values of $\alpha<0(>0)$ the term $\alpha{u}$ injects (dissipates) energy and $\beta>0$ suppresses energy buildup at large length scales. $\alpha<0$ and $\beta$ define a characteristic velocity scale, $v=\sqrt{|\alpha|/\beta}$, for polar-ordered motion. The ($\Gamma_{0},\Gamma_{2}$) term has the structure of the Swift-Hohenberg equation [42]. With $\Gamma_{0}<0$ and $\Gamma_{2}>0$, this term inherently drives the system toward an instability at the length scale $\Lambda=2\pi\sqrt{\tfrac{2\Gamma_{2}}{|\Gamma_{0}|}}$, velocity $v_{0}=\sqrt{\tfrac{|\Gamma_{0}|^{3}}{\Gamma_{2}}}$ and the corresponding instability growth timescale is $\tau_{\Gamma}=\tfrac{\Lambda}{v_{0}}$. It is the interplay between these terms that leads to chaotic, self-sustaining, turbulent-like flow patterns.
@@
Simulation details— We solve Eq. (1) using a de-aliased pseudo-spectral method with periodic boundary conditions. Simulations are performed in a two-dimensional square domain of size, L=160=160 with $4096^{2}$ grid points, ensuring adequate resolution of fine-scale features. Time integration employs the second-order integrating factor scheme (IFRK2) with a time step $\Delta t=2\times 10^{-4}$. The simulations were initialized with a random vorticity field, and in all cases, we set the parameters fixed (except $\alpha$) unless otherwise stated, $\lambda_{0}=3.5$, $\beta=0.5$, $\Gamma_{2}=9\times 10^{-5}$, $\Gamma_{0}=-4.5\times 10^{-2}$. These values are chosen based on previous studies that align with experimentally accessible regimes [16]. For more details, see section S1. To investigate the role of activity, we systematically vary the parameter $\alpha$ between -9 to -1. Using average velocities observed in experiments on B. subtilis at normal oxygen concentration, $\sim 25\mu m/s$  [48, 4, 16], our simulation parameters map to physical $v_{rms}$ in the range $4.5-70\mu m/s$, consistent with experimental observations [21], with the full correspondence detailed in  S1.
Figure 2: (a) The normalized velocity correlation length $\zeta/\Lambda$ versus $\alpha$. Inset: The spatial velocity correlation function, defined as $C_{v}(r)=\langle\mathbf{u}(\mathbf{r}_{0})\cdot\mathbf{u}(\mathbf{r}_{0}+\mathbf{r})\rangle$ is shown as a function of distance $r$ for increasing values of $|\alpha|$. $\zeta$ is extracted by fitting the decay of $C_{v}(r)\sim e^{-r/\zeta}$. For $\alpha>\alpha_{c}$, $\zeta/\Lambda\sim 1$, whereas it increases significantly for $\alpha<\alpha_{c}$, indicating the emergence of long-range velocity correlations in the high-activity regime. (b) Probability distribution function of the normalized velocity component $u_{x}/\sqrt{|\alpha|/\beta}$ for different activity values $\alpha$. A transition from a unimodal to a bimodal distribution is observed as $\alpha$ decreases beyond the critical point $\alpha_{c}$. (c) The order parameter $\phi=E_{\rm active}-E_{\Lambda}-E_{\rm kin}$ is plotted as a function of $\alpha$ for three values of $\tau_{\Gamma}$ obtained by varying $\Gamma_{2}$. The baseline case with $\tau_{\Gamma}=0.4$ ($\Gamma_{2}=9.0\times 10^{-5}$, as specified in the parameters) is shown by circles, while increased and decreased $\tau_{\Gamma}$ are represented by triangles and squares, respectively. For intermediate and larger $\Gamma_{2}$, $\phi$ remains positive below the corresponding critical activity $\alpha_{c}$ and becomes negative above it, marking the transition. In contrast, for the smallest $\Gamma_{2}$, $\phi$ stays negative across all $\alpha$, and no transition point $\alpha_{c}$ is observed.
@@
Spatial distribution of vortex centers— We show in Fig. 1 (a) the plots of vorticity field $\omega(\mathbf{x},t)=\nabla\times\mathbf{u}(\mathbf{x},t)$ for $\alpha=-1$ (I) and 9-9 (II). The field $\omega$ undergoes a structural transformation as $\alpha$ decreases, with large-scale spatial inhomogeneities becoming more prominent at lower $\alpha$ or, equivalently, higher activity. A similar feature has also been observed in earlier studies of active turbulence with increasing activity [26, 35]. These inhomogeneities are evident from the large-scale vortical structures as seen in Figs. 1(a III) and  1(a IV). In active matter systems, such large-scale spatial variations are a hallmark of giant number fluctuations, wherein the standard deviation ($\Delta n$) of the number of observables — in this case, vortex centers ($n$) — grows anomalously large and scale with the average as $\Delta n\sim\bar{n}^{\delta}$, with the exponent $\delta$ reaching values close to unity in two dimensions [3, 28, 9]. In the asymptotic limit of $\rm n$, $\delta=0.5$ for systems in equilibrium and away from continuous phase transitions. In analogy, we find giant number fluctuations of the vortex centers in active turbulence. To quantify these fluctuations, we measure the average number of vortices, $\bar{n}$, and the corresponding standard deviation, $\Delta n$, within a subregion of size $l^{2}$. As $\alpha$ decreases, $P(n/\bar{n})$ broadens, revealing stronger fluctuations in the number of vortex centers (Fig. S4). In Fig. 1(c) we plot $\Delta n$ versus $\bar{n}$ for different activities and find that $\Delta\rm n\sim\bar{n}^{\delta}$, with $\delta\simeq 0.5$ for $\alpha\geq-5$ and $\delta>0.5$ for $\alpha<-5$. The values of $\delta$ approach 1 for large activity values, see inset of Fig. 1(c). We find that giant number fluctuations persist up to a subregion of size order $20\Lambda$ $\sim\mathcal{O}$(correlation length) (Fig. 2(a)), beyond which number fluctuations start decaying toward homogeneity. Notably, the transition of $\delta$ at a critical value of the activity parameter, $\alpha_{c}=-5$, matches with the previously identified critical activity by the group of Ray [26, 27] through the onset of universality in energy spectra, increased intermittency, and other hallmarks of turbulence. Note that with increasing activity, although the vorticity at individual vortex centers increases, the total number of such centers decreases S2(b).
Giant number fluctuations are typically attributed to long-range correlations in density fluctuations [5, 45, 43, 9], seen as a divergence of the structure factor in the long-wavelength limit. However here, such density fluctuations are absent because the fluid is incompressible. Nevertheless, one can define an effective number density for vortex centers, $\rho_{l}(\bm{x},t)=n/l^{2}$. The static structure factor of $\rho_{l}$ grows with decreasing $k$ and peaks at a finite wavenumber (corresponding length is $\sim 50\Lambda$, for $\alpha=-6$), indicating the presence of giant number fluctuations at a mesoscale (see section S3), Fig S5. Consistently, the structure factor remains finite in the long-wavelength limit, reflecting the fact that these giant number fluctuations arise at a mesoscale and not at system sizes $\sim 400\Lambda$. Furthermore, when giant number fluctuations are absent, corresponding to $\alpha>-5$, the structure factor does not have a characteristic peak and saturates to a constant value as $k\to 0$ (see Fig S5), indicating homogenization of fluctuations in $\rho_{l}$ [31].
@@
The methodology for identifying vortex centers is provided in section S2. While the fluctuations in number reflect statistical deviations, a more structural perspective on the spatial inhomogeneity of vortex centers is offered by Voronoi tessellation. The results of this analysis are consistent with the emergence of giant number fluctuations at high activity levels and show a transition at $\alpha_{c}=-5$ (see section S4).
@@
Emergence of local ordering— The emergence of giant number fluctuations raises the question of what structural reorganization in the velocity field underlies this transition. To address this, we closely examine the structure of the vorticity and velocity fields. For $\alpha<\alpha_{c}$, there are a few large vortical regions of the size $\sim 25\Lambda$ characterized by low or nearly vanishing vorticity (see Fig. 1(a IV)) interspersed with vortical regions of the size $\mathcal{O}(\Lambda)$. In Fig 1(b I), we plot the velocity streamlines corresponding to these large vortical regions of vanishing vorticity and note that they are highly aligned, indicating local ordering. This contrasts sharply with other regions of the same field Fig 1(b II), where the flow remains disordered and lacks directional alignment. Furthermore, in Fig 2(a) we plot velocity correlation function $C_{v}(r)=\langle\mathbf{u}(\mathbf{r}_{0})\cdot\mathbf{u}(\mathbf{r}_{0}+\mathbf{r})\rangle$ versus $r$ and see that there is a sharp transition followed by a monotonic increase in correlation length, $\zeta$, with decreasing $\alpha$, for $\alpha<\alpha_{c}$. For sufficiently negative $\alpha$ (e.g., $\alpha\lesssim-6$), the spatial velocity correlation length reaches up to $\mathcal{O}(10\Lambda)$ and for $\alpha=-9$ we find $\zeta\simeq 25\Lambda$, consistent with the lengths of the large vortical regions seen in Fig. 1 (a II). For $\alpha>\alpha_{c}$, we note from Fig. 1 (a I) and (a III), the largest vortical structures are $\sim\mathcal{O}(\Lambda)$ and velocity streamlines are chaotic throughout domain, see Fig. S2(a). Musacchio and group [34] have reported a similar trend, however, they observe a saturation of the correlation length to an asymptotic value for high activities, rather than the monotonic growth found in our system. This saturation is attributed to the confined geometry employed in their study.
Signatures of local ordering can also be seen in the probability distribution function (PDF) of the velocity field: we plot in Fig. 2(b) the PDF of a component of the velocity field for different activities and observe that the PDF transitions from a unimodal distribution centered at $\langle u\rangle=0$ for $\alpha>\alpha_{c}$ to a bimodal distribution for $\alpha<\alpha_{c}$. The two peaks in the bimodal state progressively approach $\pm\sqrt{\frac{\left|\alpha\right|}{\beta}}$, consistent with the predictions of the Toner-Tu model for self-propelled active systems [44]. This highlights competition between active alignment-induced ordering Toner-Tu terms and spatial modulation introduced by higher-order gradient terms.
To gain a deeper understanding of the local ordering that emerges along with the large vortical structure in Fig. 1(a IV), we analyze the velocity field around the center of this vortex. We represent the velocity field by shifting the origin to the vortex center and expressing it in polar coordinates as follows:
$\mathbf{u}(r)=u_{r}(r)\,\hat{\mathbf{r}}+u_{\theta}(r)\,\hat{\bm{\theta}}$
where $u_{r}$ and $u_{\theta}$ denote the radial and azimuthal velocity components, respectively. Our numerical simulations show that $u_{r}\ll u_{\theta}$ and $u_{\theta}\sim\sqrt{|\alpha|/\beta}$ for $r>\Lambda$ (Fig. S7), consistent with the asymptotic solution [35]. To investigate the behavior near the vortex center, i.e., in the regime $r\ll\Lambda$, we assume a power-law ansatz $u_{\theta}\sim r^{a}$. In this regime, the dynamics are dominated by the linear scaling $u_{\theta}\sim r$ (see detailed discussion in Sec. S5). However, due to resolution limits, the small-$r$ behavior cannot be reliably extracted from the simulations and thus remains beyond the scope of the present work. Although this scaling may be related to the dynamics of vortex defects reported in the incompressible Toner–Tu (ITT) system [37], establishing this connection requires further study, which we leave for future investigation.
@@
The global phase behavior is not dictated by a single control parameter but instead emerges from a balance between multiple energetically distinct contributions. The first is the energy associated with activity-driven motion, quantified as $E_{\rm active}=\frac{v^{2}}{2}.$ The second arises from the destabilizing active stresses introduced by the generalized Swift–Hohenberg-like term, which generates intrinsic instabilities and injects energy into the system at intermediate length scales. The energy associated with this contribution is given by $E_{\Lambda}=\frac{|\Gamma_{0}|^{3}}{\Gamma_{2}}.$ The third component is the kinetic energy of the flow, $E_{\rm kin}(\equiv\int_{\mathbf{x}\in L^{2}}\frac{1}{2}\mathbf{u}\cdot\mathbf{u}d\mathbf{x}$). This energetic framework allows us to define an effective order parameter, $\phi$, that captures the interplay between these different forms of energy input and transfer.
@@
We hypothesize that the relative magnitudes of $E_{\rm active}$, $E_{\Lambda}$, and $E_{\rm kin}$ determine the qualitative state of the system — whether it exhibits local ordering identified by long correlation lengths or remains in a disordered phase. For $\alpha<\alpha_{c}$, the order parameter $\phi$ becomes positive, indicating that the active energy $E_{active}$ dominates over other contributions. This leads to a breakdown of homogeneity and the emergence of local ordering. In contrast, $\phi<0$ corresponds to a homogeneous, disordered state where active energy is insufficient to induce such organization, Fig. 2(c).